# Phase Diagram and Crossover Phases of Topologically Ordered Graphene Zigzag Nanoribbons: Role of Localization Effects

### 09 Jan 2024 v2

@@ 1 Introduction
Is the formation of a fractional charge [1, 2, 3, 4, 5, 6] a necessary and sufficient condition [7] for topologically ordered insulators [8] such as fractional quantum Hall systems [9] and interacting disordered zigzag graphene nanoribbons [10] (ZGNRs) with anyonic fractional charges? This issue is related to whether the electron localization effects of doped systems destroy or enhance the topological order [11, 12, 13] and quantization of fractional charges. In a Laughlin state on a sphere (with no edges present) with weak disorder, the added electrons fractionalize and create a quasi-degenerate peak within the gap of the tunneling density of states (DOS). Electron localization [14] is expected to stabilize these fractional charges of the quasi-degenerate midgap states because these localized quasi-degenerate energy states are spatially separated from each other, as explained in [15] (if fractional charges are delocalized they overlap and become ill-defined). However, excessive disorder is considered detrimental to topological order.
@@
Although weak disorder leads to formation of fractional charges, strong disorder may destroy them. Similar to fractional quantum Hall systems, the topological order of a ZGNR is not immediately destroyed upon doping because electron localization partially suppresses quantum fluctuations between quasi-degenerate mid-gap states. The system may still be an insulator with a fractional charge. However, in the presence of strong disorder or doping, zigzag edge antiferromagnetism is expected to diminish, thereby affecting the topological order. (Away from the low doping region, a disordered anyon phase with a distorted edge spin density wave was discovered [31].) These results suggest the possibility of multiple topological phase transitions in the zigzag ribbons. What is the nature of these topological phase transitions and the physical properties of the ground states? Does the presence of a fractional charge imply a universal value of the TEE? Does the TEE become nonuniversal and vary [35] with an increase in the disorder strength or doping level?
The main results of our paper are as follows: the phase diagram of ZGNRs in the parameter space comprising on-site repulsion ($U$), disorder strength ($\Gamma$), and doping concentration ($\delta N/N_{s}$) ($\delta N$ and $N_{s}$ are, respectively, the number of doped electrons and the total number of sites in the ribbon). We identified a number of different phases, including those with topological order, quasi-topological order, and no order. Each of these phases is defined by the value of TEE $\beta$ and its variance. These properties of $\beta$ are related to the presence or absence of charge fractionalization and charge transfer correlations between zigzag edges. When both of these properties are present, along with correlations leading to spin-charge separation, $\beta$ is universal with small variances. In this case, we investigated the tunneling of fractional charges under an applied voltage between the zigzag edges of undoped topologically ordered zigzag ribbons and observed a possible zero-bias anomaly. The other two types of phases fall into the category of quasi-topological order. We refer to these phases as crossover phases, in which the variance of $\beta$ is significant. In one of these phases, both $e^{-}/2$ fractional charges and spin-charge separation are absent; however, the charge transfer (±$\pm e^{-}/2$) correlations exist between the zigzag edges. Another phase may contain stable $e^{-}/2$ fractional charges but lacks charge transfer correlations between the zigzag edges. We also investigated the ground state and zigzag edge properties of the various crossover and nontopological phases.
@@ 2 Model Hamiltonian
There may be several nearly degenerate Hartree-Fock ground states. We select the Hartree-Fock initial ground state such that $\langle n_{i,\sigma}\rangle$ represents a paramagnetic state with a small spin splitting. In addition, we choose small random values for $\langle h_{ix}\rangle$ and $\langle h_{iy}\rangle$ ( these values do not significantly affect the final results). The Hartree-Fock matrix dimension scales with the number of carbon atoms, which is typically less than $50,000$. The Hartree-Fock eigenstates and eigenenergies are self-consistently computed, requiring approximately $20$ iterations. The TEE is computed using the disorder-averaging results from numerous disorder realizations. Here, we used gpu to speed up the solution of the Hartree-Fock matrix. The gpu calculations were computationally intensive and were performed on a supercomputer. In the presence of disorder and in the low-doping region, the obtained Hartree-Fock ground-state properties with solitons are in qualitative agreement with those of the density matrix renormalization group in the matrix product representation [31]. (In this work we do not investigate the high doping region. Obtaining density matrix renormalization group results in this region is challenging because the computation is rather time consuming, making it impossible to determine the true ground state among the nearly degenerate Hartree-Fock ground states.)
@@ 3 Effects of Anderson localization
Therefore, when the energy difference is sufficiently small, the states may become delocalized, and the overlap between the localized states of the left and right wells is significant. Symmetric and antisymmetric states of a quantum double well serve as good examples of these delocalized states.
Anderson localization plays a crucial role in the quantization of fractional charges [15]. The effects of Anderson localization can be described using self-consistent Hartree-Fock approximation [38, 39]. The first important effect of Anderson localization in the presence of on-site repulsion is that quasi-degenerate localized midgap states are spatially separated [14, 32], leading to well-defined fractional charges [15]. (In low-doped ZGNRs, added electrons fractionalize and form a narrow peak in the DOS near $E=0$ consisting of quasi-degenerate localized states  [31].) Figure 2(d) displays the probability densities of such two midgap states carrying fractional charges. These gap-edge states are mixed chiral gap-edge states [24], whose probability densities peak at the two edges and rapidly decays inside the ribbon. Note that these states do not overlap with each other. Non-interacting electrons of disordered ZGNRs also display mixed chiral gap-edge states near $E=0$. However, in the weak-disorder regime, the overlap between nearly degenerate states is small but not negligible. Thus well-defined fractional charges do not readily form in non-interacting disordered ZGNRs.
@@
The results of correlation length $\xi$ are shown in Figure 2. For disordered cases with $\Gamma\neq 0$, the correlation length is obtained by averaging over several disorder realizations. Anderson localization leads to a reduction in the correlation length when compared to disorder-free ribbons, as demonstrated in Figure 2(b). (Disorder-free ZGNRs exhibit a large correlation length for small $U$.) Therefore, compared to the non-disordered case, we can use a smaller Wilson loop [12] to calculate the TEE of disordered interacting ZGNRs. In contrast, doping results in an increased correlation length, as depicted in Figure 2(c).
@@ 4 Phase Diagram
In undoped ZGNRs, a topologically ordered phase is observed in regions $\Gamma/U\lesssim 1$ and $U\lesssim t$, see Figure 3(c). The topological phase transition into the symmetry-protected phase at $\Gamma=0$ is abrupt, consistent with findings in [30] (TEE of the symmetry-protected phase phase is zero). There are also other topological phase transitions, but they are smooth transitions with crossover regions. Figures 3(e)-(g) reveal the presence of crossover regimes lying beyond the topologically ordered phase with an increase in the disorder, doping, and interaction strength. The phase boundaries between topologically ordered and normal phases are “blurred,” indicating the presence of crossover phases (comprising two types, crossover phase I and crossover phase II). The numerical results of the TEE are shown in Figure 3(c). The error bars in this figure include not only random fluctuations caused by disorder but also the uncertainties arising from the extrapolation process of the finite scaling analysis. As $\Gamma/U$ increases, $\beta$ decreases (indicated by the red line in Figure 3(b)). The value of the TEE thus changes across a crossover phase. Within this phase, $\beta$ exhibits a large variance, but the average values are not zero, which implies that the topological order is not completely destroyed. In this regime, the TEE becomes nonuniversal and decays.
One can use a different but equivalent procedure to determine the phase diagram. We have verified that a similar phase diagram can be obtained by analyzing the presence of fractional charges and nonlocal correlations between the opposite zigzag edges. For each grid point $(\Gamma,U,\delta N)$ in the parameter space, we find the ground state and investigate whether the gap-edge states display fractional charges and whether nonlocal correlations exist between the opposite zigzag edges. By ultilizing this method, we have successfully recovered the phase diagram depicted in Figure 3(d).
@@ 7 Crossover phase II
For undoped ZGNRs, another crossover phase emerges when $\Gamma\gg U$ but $U/t\lesssim 1$. We call this phase crossover II where the disorder strength $\Gamma$ is the dominant energy. As $\Gamma$ increases, the topologically ordered phase undergoes a gradual transition into the crossover phase II, as depicted in Figure 3(e). Concurrently, the gap is progressively filled with states, as illustrated in the upper graph of Figure 8(a). Similar to the crossover I phase (Figure 7(a)), $\beta$ is finite with a significant variance. However, in this phase, there are no charge-transfer correlations between the zigzag edges, as evident from the lower graph in Figure 8(a). However, some fractional charges may exist, as shown in Figure 8(b). This observation aligns with the following obtained results: (i) Some changes in the edge occupation number are $\delta n_{i,\sigma}\approx\pm 1/2$. (ii) There are gap-edge states with $q_{A}\approx 1/2$. (Here $q_{A}=\sum_{i\in A}|\psi_{i\sigma}(E)|^{2}$, where $\psi_{i\sigma}(E)$ is the Hartree-Fock eigenstate with energy $E$, see [26]. The probability densities are summed over all sites of the A sublattice.) However, although the mean value of $q_{A}$ is $1/2$, its variance in the energy interval $[E-\delta E,E+\delta E]$ is large because $q_{A}$ varies substantially within this interval, as shown in Figure 8(a). Despite this, a fractional charge of a state in the interval $[E-\delta E,E+\delta E]$ near $E=0$ does not overlap significantly with the probability densities of other fractional and non-fractional states in the same energy interval, provided that $\delta E$ is small (for $U\sim t$ and $\Gamma\sim t$, this happens when $\delta E\sim 0.01t$). (However, in the absence of on-site repulsion, multiple states with the same value of $\delta E$ overlap.) We posit that the interplay between localization and on-site repulsion accounts for these non-overlapping states, as previously discussed at the outset of Sec.3. Nevertheless, the impact of quantum fluctuations[15] must not be disregarded, as significant overlap with states in the adjacent energy intervals may occur.
@@
So far, our investigation has primarily focused on the undoped case. Upon doping, the disorder-free ZGNRs exhibit edge spin density waves instead of edge ferromagnetism observed in undoped ribbons. If disorder is added to a doped ZGNR, the spin waves become distorted [31]: There is a topological phase transition from modulated ferromagnetic edges at zero doping to distorted spin-wave edges at finite doping. Our results indicated that, under substantial doping, this phase is also a crossover phase II. The dependence of the mean value of $q_{A}$ of the states within the midgap peak on the number of doped electrons is illustrated in Figure 8(d) (the DOS shows a sharp peak at the midgap energy as seen in Figure 8(c)). At a low doping concentration, the disorder-averaged value of $q_{A}$ of the midgap peak is close to $0.5$, and the midgap states display well-defined fractional charges with small variance, as discussed below Fig. 2 (note that the width of the midgap peak is $\delta E\sim 0.005t$). As the doping concentration increases further, $q_{A}$ significantly deviates from 0.5, and simultaneously, the DOS midgap peak starts to decrease [31]. These findings imply that even though fractional charges can still be observed, their number decreases with an increase in doping. The gradual change in $q_{A}$ as a function of $\delta N/N_{\textrm{s}}$ indicates that the transition from the phase of the distorted ferromagnetic edge to the phase of the distorted edge spin-wave is not sharp. Figure 8(e) shows how $\beta$ decreased with an increase in $\delta N/N_{\textrm{s}}$. For a large $\delta N/N_{\textrm{s}}$, it is computationally demanding to calculate $\beta$ due to the anticipated longer correlation length, as observed in Figure 2(c)).
@@ 8 Strongly disordered and strongly repulsive phases
Figure 9: (a) Site spins and site occupation numbers for spin-up and -down are shown for a ribbon size $(L,W)=(100,64)$. The calculations were performed for two different sets of parameters: the left one with $(U,\Gamma,\delta N)=(t,0.03t,0)$ and the right other with $(U,\Gamma,\delta N)=(t,15t,0)$. Both cases were computed by single disorder realization $N_{D}=1$. (b) Edge occupation numbers for spin-up, spin-down, and edge site spins in strongly disordered regime with $(U,\Gamma,\delta N)=(t,15t,0)$. The geometry of the ribbon is $(L,W)=(100,8)$ and these quantities are obtained by single disorder realization $N_{D}=1$. (c) $E$ $q_{A}$ graph with $(L,W)=(100,8)$, $(U,\Gamma,\delta N)=(7t,0.5t,0)$ ($N_{D}=1$). Red (blue) dots are states of spin up (down). Solid red and blue lines represent $q_{A}$ values of spin-up and -down states for $\Gamma=0$. These states correspond to zigzag edge states localized either on the upper or lower zigzag edges. Disorder-free DOS for $(U,\Gamma,\delta N)=(7t,0,0)$ shown in the inset has a large gap, $\Delta\approx 6t$. (d) Probability density of several states near gap edges ±$\pm\Delta/2$ (horizontal thin black lines) with $q_{A}\approx 0.5$ for single disorder realization $N_{D}=1$ and $(U,\Gamma,\delta N)=(7t,0.5t,0)$. The ribbon’s dimensions are set to $(L,W)=(100,8)$. (e) DOS for a ribbon of dimensions $(L,W)=(100,64)$ with $(U,\Gamma,\delta N)=(4t,4t,0)$ ($N_{D}=100$ realizations were made). The computed TEE value is $\beta=0$. (f) Zigzag edge occupation numbers $n_{i,\sigma}$, $n_{i}=n_{i,\uparrow}+n_{i,\downarrow}$ and site spins $S_{z}$ in the strong repulsive regime with parameters $(U,\Gamma,\delta N)=(5t,0.01t,0)$ with a ribbon’s dimension of $(L,W)=(120,8)$. These results are based on a single disorder realization ($N_{D}=1$). (g) The same quantities are computed for the parameters $(U,\Gamma,\delta N)=(5t,0.1t,0)$ using a ribbon with the dimensions mentioned in (f). At $\Gamma=0.1t$, more sites exhibit total site occupation numbers $n_{i}\approx 2$ in comparison to $\Gamma=0.01t$.

### 05 Apr 2024 v3

@@ 1 Introduction
Is the formation of a fractional charge [1, 2, 3, 4, 5, 6] a necessary and sufficient condition [7] for topologically ordered insulators [8] such as fractional quantum Hall systems [9] and interacting disordered zigzag graphene nanoribbons [10] (ZGNRs) with anyonic fractional charges? This issue is related to whether the electron localization effects destroy or enhance the topological order [11, 12, 13] and quantization of fractional charges. In a Laughlin state on a sphere (no edges present) and under weak disorder, the added electrons fractionalize and create a quasi-degenerate peak within the gap of the tunneling density of states (DOS). Electron localization [14] is expected to stabilize these fractional charges of the quasi-degenerate midgap states because these localized quasi-degenerate energy states are spatially separated from each other, as explained in [15] (if fractional charges are delocalized they overlap and become ill-defined). However, excessive disorder is considered detrimental to topological order.
@@
Although weak disorder leads to formation of fractional charges, strong disorder may destroy them. Similar to fractional quantum Hall systems, the topological order of a ZGNR is not immediately destroyed upon doping because electron localization partially suppresses quantum fluctuations between quasi-degenerate mid-gap states. The system may still be an insulator with a fractional charge. However, in the presence of strong disorder or doping, zigzag edge antiferromagnetism is expected to diminish, thereby affecting the topological order. (In the doped region, a disordered anyon phase with a distorted edge spin density wave was discovered [31].) These results suggest the possibility of multiple topological phase transitions in the zigzag ribbons. What is the nature of these topological phase transitions and the physical properties of the ground states? Does the presence of a fractional charge imply a universal value of the TEE? Does the TEE become nonuniversal and vary [35] with an increase in the disorder strength or doping level?
The main results of our paper are as follows: the phase diagram of ZGNRs is found in the parameter space comprising on-site repulsion ($U$), disorder strength ($\Gamma$), and doping concentration ($\delta N/N_{s}$) ($\delta N$ and $N_{s}$ are, respectively, the number of doped electrons and the total number of sites in the ribbon). We identified a number of different phases, including those with topological order, quasi-topological order, and no order. Each of these phases is defined by the value of TEE $\beta$ and its variance. These properties of $\beta$ are related to the presence or absence of charge fractionalization and charge transfer correlations between zigzag edges. When both of these properties are present, along with correlations leading to spin-charge separation, $\beta$ is universal with small variances. In this case, we investigated the tunneling of fractional charges under an applied voltage between the zigzag edges of undoped topologically ordered zigzag ribbons and observed a possible zero-bias anomaly. The other two types of phases fall into the category of quasi-topological order. We refer to these phases as crossover phases, in which the variance of $\beta$ is significant. In one of these phases, both $e^{-}/2$ fractional charges and spin-charge separation are absent; however, the charge transfer (±$\pm e^{-}/2$) correlations exist between the zigzag edges. Another phase may contain weakly stable $e^{-}/2$ fractional charges but lacks charge transfer correlations between the zigzag edges. We also investigated the ground state and zigzag edge properties of the various crossover and nontopological phases.
@@ 2 Model Hamiltonian
There may be several nearly degenerate Hartree-Fock ground states. We select the Hartree-Fock initial ground state such that $\langle n_{i,\sigma}\rangle$ represents a paramagnetic state with a small spin splitting. In addition, we choose small random values for $\langle h_{ix}\rangle$ and $\langle h_{iy}\rangle$ (these values do not significantly affect the final results). The Hartree-Fock matrix dimension scales with the number of carbon atoms, which is typically less than $50,000$. The Hartree-Fock eigenstates and eigenenergies are self-consistently computed, requiring approximately $20$ iterations. The TEE is computed using the disorder-averaging results from numerous disorder realizations. Here, we used GPU to speed up the solution of the Hartree-Fock matrix. The GPU calculations were computationally intensive and were performed on a supercomputer. In the presence of disorder and in the low-doping region, the obtained Hartree-Fock ground-state properties with solitons are in qualitative agreement with those of the density matrix renormalization group in the matrix product representation [31]. However, quantum fluctuations are not included in the Hartree-Fock calculation, so it may overestimate the stability of fractional charges. (In this work we do not investigate the high doping region. Obtaining density matrix renormalization group results in this region is challenging because the computation is rather time consuming, making it impossible to determine the true ground state among the nearly degenerate Hartree-Fock ground states.)
@@ 3 Effects of Anderson localization
Therefore, when the energy difference is sufficiently small, the states may become delocalized, and the overlap between the localized states is significant. Symmetric and antisymmetric states of a quantum double well serve as good examples of these delocalized states.
Anderson localization plays a crucial role in the quantization of fractional charges [15]. The effects of Anderson localization can be described using self-consistent Hartree-Fock approximation [38, 39]. The first important effect of Anderson localization in undoped ribbons in the presence of on-site repulsion is that quasi-degenerate localized midgap states are spatially separated [14, 32], leading to well-defined fractional charges [15]. (In low-doped ZGNRs, added electrons fractionalize and form a narrow peak in the DOS near $E=0$ consisting of quasi-degenerate localized states  [31].) Figure 2(d) displays the probability densities of such two midgap states carrying fractional charges. These gap-edge states are mixed chiral gap-edge states [24], whose probability densities peak at the two edges and rapidly decays inside the ribbon. Note that these states do not overlap with each other. Non-interacting electrons of disordered ZGNRs also display mixed chiral gap-edge states near $E=0$. However, in the weak-disorder regime, the overlap between nearly degenerate states is small but not negligible. Thus well-defined fractional charges do not readily form in non-interacting disordered ZGNRs.
@@
The results of correlation length $\xi$ are shown in Figure 2. For disordered cases with $\Gamma\neq 0$, the correlation length is obtained by averaging over several disorder realizations. Anderson localization leads to a reduction in the correlation length when compared to disorder-free ribbons, as demonstrated in Figure 2(b). (Disorder-free ZGNRs exhibit a large correlation length for small $U$.) Therefore, compared to the non-disordered case, we can use a smaller Wilson loop [12] to calculate the TEE of disordered interacting ZGNRs. In contrast, doping results in an increased correlation length, as depicted in Figure 2(c). We stress that, in computing the correct TEE, it is important that the correlation length is small in comparison to the dimensions of the Wilson loop.
@@ 4 Phase Diagram
In undoped ZGNRs, a topologically ordered phase is observed in regions $\Gamma/U\lesssim 1$ and $U\lesssim t$, see Figure 3(c). The topological phase transition into the symmetry-protected phase at $\Gamma=0$ is abrupt, consistent with findings in [30] (TEE of the symmetry-protected phase phase is zero). There are also other topological phase transitions, but they are smooth transitions with crossover regions. Note that the system is not topologically ordered at $U=0$: our Hartree-Fock results show that fractional charges overlap with each other and charge-transfer correlations are absent. Figures 3(e)-(g) reveal the presence of crossover regimes lying beyond the topologically ordered phase with an increase in the disorder, doping, and interaction strength. The phase boundaries between topologically ordered and normal phases are “blurred,” indicating the presence of crossover phases (comprising two types, crossover phase I and crossover phase II). The numerical results of the TEE are shown in Figure 3(c). The error bars in this figure include not only random fluctuations caused by disorder but also the uncertainties arising from the extrapolation process of the finite scaling analysis. As $\Gamma/U$ increases, $\beta$ decreases (indicated by the red line in Figure 3(b)). The value of the TEE thus changes across a crossover phase. Within this phase, $\beta$ exhibits a large variance, but the average values are not zero, which implies that the topological order is not completely destroyed. In this regime, the TEE becomes nonuniversal and decays.
One can use a different but equivalent procedure to determine the phase diagram. We have verified that a similar phase diagram can be obtained by analyzing the presence of fractional charges and nonlocal correlations between the opposite zigzag edges. For each grid point $(\Gamma,U,\delta N)$ in the parameter space, we find the ground state and investigate whether the gap-edge states display fractional charges and whether nonlocal correlations exist between the opposite zigzag edges. By utilizing this method, we have successfully recovered the phase diagram depicted in Figure 3(d).
@@ 7 Crossover phase II
For undoped ZGNRs, another crossover phase emerges when $\Gamma\gg U$ but $U/t\lesssim 1$. We call this phase crossover II where the disorder strength $\Gamma$ is the dominant energy. As $\Gamma$ increases, the topologically ordered phase undergoes a gradual transition into the crossover phase II, as depicted in Figure 3(e). Concurrently, the gap is progressively filled with states, as illustrated in the upper graph of Figure 8(a). Similar to the crossover I phase (Figure 7(a)), $\beta$ is finite with a significant variance. However, in this phase, there are no charge-transfer correlations between the zigzag edges, as evident from the lower graph in Figure 8(a). Some fractional charges may exist, as shown in Figure 8(b), but they are expected to be subject to greater quantum fluctuations.
This observation aligns with the following obtained results: (i) Some changes in the edge occupation number are $\delta n_{i,\sigma}\approx\pm 1/2$. (ii) There are gap-edge states with $q_{A}\approx 1/2$. (Here $q_{A}=\sum_{i\in A}|\psi_{i\sigma}(E)|^{2}$, where $\psi_{i\sigma}(E)$ is the Hartree-Fock eigenstate with energy $E$, see [26]. The probability densities are summed over all sites of the A sublattice.) However, although the mean value of $q_{A}$ is $1/2$, its variance in the energy interval $[E-\delta E,E+\delta E]$ is large because $q_{A}$ varies substantially within this interval, as shown in Figure 8(a). Despite this, a fractional charge of a state in the interval $[E-\delta E,E+\delta E]$ near $E=0$ does not overlap significantly with the probability densities of other fractional and non-fractional states in the same energy interval, provided that $\delta E$ is small (for $U\sim t$ and $\Gamma\sim t$, this happens when $\delta E\sim 0.01t$). (However, in the absence of on-site repulsion, multiple states with the same value of $\delta E$ overlap.) We posit that the interplay between localization and on-site repulsion accounts for these non-overlapping states, as previously discussed at the outset of Sec.3. Nevertheless, the impact of quantum fluctuations [15] must not be disregarded, as significant overlap with states in the adjacent energy intervals may occur. (The Hartree-Fock calculation, which ignores quantum fluctuations, may thus overestimate the stability of fractional charges.)
@@
So far, our investigation has primarily focused on the undoped case. Upon doping, the disorder-free ZGNRs exhibit edge spin density waves, with the opposite edges still being antiferromagnetically coupled. This is in contrast to the edge ferromagnetism observed in undoped ribbons. If disorder is added to a doped ZGNR, the spin waves become distorted [31]: There is a topological phase transition from modulated ferromagnetic edges at zero doping to distorted spin-wave edges at finite doping. Our results indicated that, under substantial doping, this phase is also a crossover phase II. The dependence of the mean value of $q_{A}$ of the states within the midgap peak on the number of doped electrons is illustrated in Figure 8(d) (the DOS shows a sharp peak at the midgap energy as seen in Figure 8(c)). At a low doping concentration, the disorder-averaged value of $q_{A}$ of the midgap peak is close to $0.5$, and the midgap states display well-defined fractional charges with small variance, as discussed below Fig. 2 (note that the width of the midgap peak is $\delta E\sim 0.005t$). As the doping concentration increases further, $q_{A}$ significantly deviates from 0.5, and simultaneously, the DOS midgap peak starts to decrease [31]. These findings imply that even though fractional charges can still be observed, their number decreases with an increase in doping. The gradual change in $q_{A}$ as a function of $\delta N/N_{\textrm{s}}$ indicates that the transition from the phase of the distorted ferromagnetic edge to the phase of the distorted edge spin-wave is not sharp. Figure 8(e) shows how $\beta$ decreased with an increase in $\delta N/N_{\textrm{s}}$. For a large $\delta N/N_{\textrm{s}}$, it is computationally demanding to calculate $\beta$ due to the anticipated longer correlation length, as observed in Figure 2(c).
@@ 8 Strongly disordered and strongly repulsive phases
Figure 9: (a) Site occupation numbers for spin-up and -down, along with the total occupation numbers, are shown for a ribbon size $(L,W)=(100,64)$. The calculations were performed for two different sets of parameters: the left one with $(U,\Gamma,\delta N)=(t,0.03t,0)$ and the right other with $(U,\Gamma,\delta N)=(t,15t,0)$. Both cases were computed by single disorder realization $N_{D}=1$. (b) Edge occupation numbers for spin-up, spin-down, and edge site spins in strongly disordered regime with $(U,\Gamma,\delta N)=(t,15t,0)$. The geometry of the ribbon is $(L,W)=(100,8)$ and these quantities are obtained by single disorder realization $N_{D}=1$. (c) $E$ $q_{A}$ graph with $(L,W)=(100,8)$, $(U,\Gamma,\delta N)=(7t,0.5t,0)$ ($N_{D}=1$). Red (blue) dots are states of spin up (down). Solid red and blue lines represent $q_{A}$ values of spin-up and -down states for $\Gamma=0$. These states correspond to zigzag edge states localized either on the upper or lower zigzag edges. Disorder-free DOS for $(U,\Gamma,\delta N)=(7t,0,0)$ shown in the inset has a large gap, $\Delta\approx 6t$. (d) Probability density of several states near gap edges ±$\pm\Delta/2$ (horizontal thin black lines) with $q_{A}\approx 0.5$ for single disorder realization $N_{D}=1$ and $(U,\Gamma,\delta N)=(7t,0.5t,0)$. The ribbon’s dimensions are set to $(L,W)=(100,8)$. (e) DOS for a ribbon of dimensions $(L,W)=(100,64)$ with $(U,\Gamma,\delta N)=(4t,4t,0)$ ($N_{D}=100$ realizations were made). The computed TEE value is $\beta=0$. (f) Zigzag edge occupation numbers $n_{i,\sigma}$, $n_{i}=n_{i,\uparrow}+n_{i,\downarrow}$ and site spins $S_{z}$ in the strong repulsive regime with parameters $(U,\Gamma,\delta N)=(5t,0.01t,0)$ with a ribbon’s dimension of $(L,W)=(120,8)$. These results are based on a single disorder realization ($N_{D}=1$). (g) The same quantities are computed for the parameters $(U,\Gamma,\delta N)=(5t,0.1t,0)$ using a ribbon with the dimensions mentioned in (f). At $\Gamma=0.1t$, more sites exhibit total site occupation numbers $n_{i}\approx 2$ in comparison to $\Gamma=0.01t$.